# The N3LO Twist-2 Matching of Helicity TMDs and SIDIS $q_{\ast}$ Spectrum

### v3

@@ 4.4.2 Time-like results in $\overline{\text{MS}}$}
$\Delta P_{qg}^{T,\text{s}(2)}=N_{f}^{3}\bigg[\left(-\frac{8}{3}\right)\zeta_{2}+\left(-\frac{8}{3}\right)H_{0,0}+\left(-\frac{8}{3}\right)H_{1,1}+\left(-\frac{8}{9}\right)H_{1}+\frac{8}{9}H_{0}+\frac{8}{3}H_{0,1}+\frac{8}{3}H_{1,0}+z\left(\left(-\frac{16}{3}\right)H_{0,1}+\left(-\frac{16}{3}\right)H_{1,0}+\left(-\frac{32}{9}\right)H_{1}+\frac{32}{9}H_{0}+\frac{16}{3}\zeta_{2}+\frac{16}{3}H_{0,0}+\frac{16}{3}H_{1,1}-\frac{16}{3}\right)+\frac{40}{9}\bigg]+C_{F}N_{f}^{2}\bigg[\left(16\zeta_{2}+16H_{-1,0}-16H_{0,0}\right)z^{2}+\left(\left(-\frac{128}{3}\right)H_{0,0,1}+\left(-\frac{112}{3}\right)H_{1,1,0}+\left(-\frac{280}{9}\right)H_{0,1}+\left(-\frac{260}{9}\right)H_{0,0}+\left(-\frac{80}{3}\right)H_{1,0,1}+\left(-\frac{80}{3}\right)H_{1,1,1}+\left(-\frac{680}{27}\right)H_{1}+\frac{280}{9}\zeta_{2}+\frac{328}{9}H_{1,1}+\frac{128}{3}H_{0,1,1}+\frac{160}{3}H_{0,1,0}+\frac{232}{3}H_{1,0}+\frac{320}{3}H_{0,0,0}+\frac{5090}{27}H_{0}+\frac{608}{3}\zeta_{3}+\frac{64\zeta_{2}}{3}H_{0}-16H_{1}\zeta_{2}-64H_{-1,0}+64H_{0,-1,0}+64H_{1,0,0}-64H_{0,0,0,0}-\frac{1157}{3}\right)z$ (139)
$\left(-\frac{304}{3}\right)\zeta_{3}+\left(-\frac{848}{9}\right)\zeta_{2}+\left(-\frac{80}{3}\right)H_{0,1,0}+\left(-\frac{64}{3}\right)H_{0,1,1}+\left(-\frac{56}{3}\right)H_{1,0}+\left(-\frac{32}{9}\right)H_{1,1}+\frac{184}{27}H_{1}+\frac{40}{3}H_{1,0,1}+\frac{40}{3}H_{1,1,1}+\frac{56}{3}H_{1,1,0}+\frac{64}{3}H_{0,0,1}+\frac{272}{9}H_{0,1}+\frac{898}{9}H_{0,0}+\frac{320}{3}H_{0,0,0}+\frac{3830}{27}H_{0}+\left(-\frac{128\zeta_{2}}{3}\right)H_{0}+8H_{1}\zeta_{2}-80H_{-1,0}-32H_{0,-1,0}-32H_{1,0,0}+32H_{0,0,0,0}+\frac{6587}{18}\bigg]+N_{f}C_{A}^{2}\bigg[\left(72H_{-1}\zeta_{2}-96H_{0}\zeta_{2}-24H_{1}\zeta_{2}+24\zeta_{2}-120\zeta_{3}+24H_{-1,0}-24H_{0,0}+48H_{-1,-1,0}-48H_{-1,0,0}-48H_{-1,0,1}-48H_{0,-1,0}+48H_{0,0,0}+48H_{0,0,1}\right)z^{2}+\left(\left(-\frac{8836}{9}\right)H_{0,1}+\left(-\frac{976}{3}\right)H_{0,0,0}+\left(-\frac{848}{3}\right)H_{-1,-1,0}+\left(-\frac{2348}{9}\right)H_{0,0}+\left(-\frac{592}{3}\right)H_{1,0,1}+\left(-\frac{592}{3}\right)H_{1,1,0}+\left(-\frac{440}{3}\right)H_{1,1,1}+\left(-\frac{3796}{27}\right)H_{0}+\left(-\frac{404}{3}\right)H_{1,0,0}+\left(-\frac{320}{3}\right)H_{-1,0,1}+\left(-\frac{308}{3}\right)\zeta_{2}+\left(-\frac{304}{3}\right)H_{-1,0}+\frac{284}{3}H_{0,0,1}+\frac{536}{3}H_{-1,0,0}+\frac{5464}{27}H_{1}+\frac{1828}{9}H_{1,1}+\frac{2068}{9}H_{1,0}+\frac{824}{3}H_{0,-1,0}+\left(-\frac{104\zeta_{2}}{3}\right)H_{-1}+\frac{472\zeta_{2}}{3}H_{1}+60H_{0}\zeta_{2}+272H_{-1}\zeta_{3}+32H_{0}\zeta_{3}+496H_{1}\zeta_{3}-184\zeta_{3}+170\zeta_{4}-416\zeta_{2}H_{-1,-1}+224\zeta_{2}H_{-1,0}+144\zeta_{2}H_{0,-1}-56\zeta_{2}H_{0,0}+336\zeta_{2}H_{0,1}+96\zeta_{2}H_{1,0}-32\zeta_{2}H_{1,1}-24H_{0,1,0}+192H_{0,1,1}-192H_{-1,-1,-1,0}+128H_{-1,-1,0,0}+320H_{-1,-1,0,1}+224H_{-1,0,-1,0}+16H_{-1,0,0,0}-128H_{-1,0,1,0}-64H_{-1,0,1,1}+32H_{0,-1,-1,0}-208H_{0,-1,0,0}-128H_{0,-1,0,1}-112H_{0,0,-1,0}+128H_{0,0,0,0}-200H_{0,0,0,1}+128H_{0,0,1,0}+352H_{0,0,1,1}+184H_{0,1,0,0}-160H_{0,1,1,1}-32H_{1,0,-1,0}-240H_{1,0,0,0}-256H_{1,0,0,1}-64H_{1,0,1,1}-320H_{1,1,0,0}-64H_{1,1,0,1}-64H_{1,1,1,0}+64H_{1,1,1,1}+\frac{5408}{3}\right)z+\left(-\frac{19762}{27}\right)H_{0}+\left(-\frac{9620}{27}\right)H_{1}+\left(-\frac{928}{3}\right)H_{0,0,0}+\left(-\frac{640}{3}\right)H_{-1,-1,0}+\left(-\frac{1502}{9}\right)H_{0,0}+\left(-\frac{352}{3}\right)H_{-1,0,1}+\left(-\frac{566}{9}\right)H_{1,1}+\left(-\frac{434}{9}\right)H_{1,0}+\frac{220}{3}H_{1,1,1}+\frac{232}{3}H_{-1,0}+\frac{304}{3}\zeta_{3}+\frac{680}{3}H_{1,0,1}+\frac{680}{3}H_{1,1,0}+\frac{776}{3}H_{0,-1,0}+\frac{2390}{9}H_{0,1}+\frac{856}{3}H_{-1,0,0}+\frac{1240}{3}H_{1,0,0}+\frac{1342}{3}\zeta_{2}+\left(-\frac{224\zeta_{2}}{3}\right)H_{1}+\frac{32\zeta_{2}}{3}H_{-1}+\frac{856\zeta_{2}}{3}H_{0}+136H_{-1}\zeta_{3}+480H_{0}\zeta_{3}-248H_{1}\zeta_{3}+63\zeta_{4}-208\zeta_{2}H_{-1,-1}+112\zeta_{2}H_{-1,0}-88\zeta_{2}H_{0,-1}+316\zeta_{2}H_{0,0}-216\zeta_{2}H_{0,1}-48\zeta_{2}H_{1,0}+16\zeta_{2}H_{1,1}-40H_{0,0,1}+4H_{0,1,0}-56H_{0,1,1}-96H_{-1,-1,-1,0}+64H_{-1,-1,0,0}+160H_{-1,-1,0,1}+112H_{-1,0,-1,0}+8H_{-1,0,0,0}-64H_{-1,0,1,0}-32H_{-1,0,1,1}-176H_{0,-1,-1,0}+216H_{0,-1,0,0}+248H_{0,0,-1,0}-632H_{0,0,0,0}-388H_{0,0,0,1}-48H_{0,0,1,0}-96H_{0,0,1,1}+204H_{0,1,0,0}+144H_{0,1,0,1}+144H_{0,1,1,0}+128H_{0,1,1,1}+16H_{1,0,-1,0}+120H_{1,0,0,0}+128H_{1,0,0,1}+32H_{1,0,1,1}+160H_{1,1,0,0}+32H_{1,1,0,1}+32H_{1,1,1,0}-32H_{1,1,1,1}-\frac{4670}{3}\bigg]+N_{f}C_{F}^{2}\bigg[\left(48\zeta_{2}+48H_{-1,0}-48H_{0,0}\right)z^{2}+\left(-80\zeta_{2}H_{0}+80\zeta_{3}H_{0}+289H_{0}+600H_{1}-64H_{-1}\zeta_{2}-176H_{1}\zeta_{2}+612\zeta_{2}+992H_{1}\zeta_{3}-368\zeta_{3}+280\zeta_{4}-192\zeta_{2}H_{-1,0}+304H_{-1,0}+128\zeta_{2}H_{0,-1}-80\zeta_{2}H_{0,0}-350H_{0,0}+208\zeta_{2}H_{0,1}-340H_{0,1}-224\zeta_{2}H_{1,0}+592H_{1,0}-32\zeta_{2}H_{1,1}+32H_{1,1}-128H_{-1,-1,0}+352H_{-1,0,0}+56H_{0,0,0}+16H_{0,0,1}-16H_{0,1,0}+32H_{0,1,1}-64H_{1,0,0}-48H_{1,0,1}-144H_{1,1,0}-48H_{1,1,1}-384H_{-1,-1,0,0}-192H_{-1,0,-1,0}+352H_{-1,0,0,0}+256H_{0,-1,-1,0}+64H_{0,-1,0,0}-192H_{0,0,-1,0}+32H_{0,0,0,0}+16H_{0,0,0,1}+192H_{0,0,1,0}+128H_{0,0,1,1}-352H_{0,1,0,0}-208H_{0,1,0,1}-240H_{0,1,1,0}-112H_{0,1,1,1}-192H_{1,0,-1,0}+672H_{1,0,0,0}+160H_{1,0,0,1}+608H_{1,0,1,0}+128H_{1,0,1,1}+32H_{1,1,0,0}+160H_{1,1,0,1}+96H_{1,1,1,0}+64H_{1,1,1,1}+\frac{3747}{2}\right)z$ (140)
[…]

### v4

@@ 4.4.2 Time-like results in $\overline{\text{MS}}$}
$\Delta P_{qg}^{T,\text{s}(2)}=N_{f}^{3}\bigg[\left(-\frac{8}{3}\right)\zeta_{2}+\left(-\frac{8}{3}\right)H_{0,0}+\left(-\frac{8}{3}\right)H_{1,1}+\left(-\frac{8}{9}\right)H_{1}+\frac{8}{9}H_{0}+\frac{8}{3}H_{0,1}+\frac{8}{3}H_{1,0}+z\left(\left(-\frac{16}{3}\right)H_{0,1}+\left(-\frac{16}{3}\right)H_{1,0}+\left(-\frac{32}{9}\right)H_{1}+\frac{32}{9}H_{0}+\frac{16}{3}\zeta_{2}+\frac{16}{3}H_{0,0}+\frac{16}{3}H_{1,1}-\frac{16}{3}\right)+\frac{40}{9}\bigg]+C_{F}N_{f}^{2}\bigg[\left(16\zeta_{2}+16H_{-1,0}-16H_{0,0}\right)z^{2}+\left(\left(-\frac{128}{3}\right)H_{0,0,1}+\left(-\frac{112}{3}\right)H_{1,1,0}+\left(-\frac{280}{9}\right)H_{0,1}+\left(-\frac{260}{9}\right)H_{0,0}+\left(-\frac{80}{3}\right)H_{1,0,1}+\left(-\frac{80}{3}\right)H_{1,1,1}+\left(-\frac{680}{27}\right)H_{1}+\frac{280}{9}\zeta_{2}+\frac{328}{9}H_{1,1}+\frac{128}{3}H_{0,1,1}+\frac{160}{3}H_{0,1,0}+\frac{232}{3}H_{1,0}+\frac{320}{3}H_{0,0,0}+\frac{5090}{27}H_{0}+\frac{608}{3}\zeta_{3}+\frac{64\zeta_{2}}{3}H_{0}-16H_{1}\zeta_{2}-64H_{-1,0}+64H_{0,-1,0}+64H_{1,0,0}-64H_{0,0,0,0}-\frac{1157}{3}\right)z$
$\left(-\frac{304}{3}\right)\zeta_{3}+\left(-\frac{848}{9}\right)\zeta_{2}+\left(-\frac{80}{3}\right)H_{0,1,0}+\left(-\frac{64}{3}\right)H_{0,1,1}+\left(-\frac{56}{3}\right)H_{1,0}+\left(-\frac{32}{9}\right)H_{1,1}+\frac{184}{27}H_{1}+\frac{40}{3}H_{1,0,1}+\frac{40}{3}H_{1,1,1}+\frac{56}{3}H_{1,1,0}+\frac{64}{3}H_{0,0,1}+\frac{272}{9}H_{0,1}+\frac{898}{9}H_{0,0}+\frac{320}{3}H_{0,0,0}+\frac{3830}{27}H_{0}+\left(-\frac{128\zeta_{2}}{3}\right)H_{0}+8H_{1}\zeta_{2}-80H_{-1,0}-32H_{0,-1,0}-32H_{1,0,0}+32H_{0,0,0,0}+\frac{6587}{18}\bigg]+N_{f}C_{A}^{2}\bigg[\left(72H_{-1}\zeta_{2}-96H_{0}\zeta_{2}-24H_{1}\zeta_{2}+24\zeta_{2}-120\zeta_{3}+24H_{-1,0}-24H_{0,0}+48H_{-1,-1,0}-48H_{-1,0,0}-48H_{-1,0,1}-48H_{0,-1,0}+48H_{0,0,0}+48H_{0,0,1}\right)z^{2}+\left(\left(-\frac{8836}{9}\right)H_{0,1}+\left(-\frac{976}{3}\right)H_{0,0,0}+\left(-\frac{848}{3}\right)H_{-1,-1,0}+\left(-\frac{2348}{9}\right)H_{0,0}+\left(-\frac{592}{3}\right)H_{1,0,1}+\left(-\frac{592}{3}\right)H_{1,1,0}+\left(-\frac{440}{3}\right)H_{1,1,1}+\left(-\frac{3796}{27}\right)H_{0}+\left(-\frac{404}{3}\right)H_{1,0,0}+\left(-\frac{320}{3}\right)H_{-1,0,1}+\left(-\frac{308}{3}\right)\zeta_{2}+\left(-\frac{304}{3}\right)H_{-1,0}+\frac{284}{3}H_{0,0,1}+\frac{536}{3}H_{-1,0,0}+\frac{5464}{27}H_{1}+\frac{1828}{9}H_{1,1}+\frac{2068}{9}H_{1,0}+\frac{824}{3}H_{0,-1,0}+\left(-\frac{104\zeta_{2}}{3}\right)H_{-1}+\frac{472\zeta_{2}}{3}H_{1}+60H_{0}\zeta_{2}+272H_{-1}\zeta_{3}+32H_{0}\zeta_{3}+496H_{1}\zeta_{3}-184\zeta_{3}+170\zeta_{4}-416\zeta_{2}H_{-1,-1}+224\zeta_{2}H_{-1,0}+144\zeta_{2}H_{0,-1}-56\zeta_{2}H_{0,0}+336\zeta_{2}H_{0,1}+96\zeta_{2}H_{1,0}-32\zeta_{2}H_{1,1}-24H_{0,1,0}+192H_{0,1,1}-192H_{-1,-1,-1,0}+128H_{-1,-1,0,0}+320H_{-1,-1,0,1}+224H_{-1,0,-1,0}+16H_{-1,0,0,0}-128H_{-1,0,1,0}-64H_{-1,0,1,1}+32H_{0,-1,-1,0}-208H_{0,-1,0,0}-128H_{0,-1,0,1}-112H_{0,0,-1,0}+128H_{0,0,0,0}-200H_{0,0,0,1}+128H_{0,0,1,0}+352H_{0,0,1,1}+184H_{0,1,0,0}-160H_{0,1,1,1}-32H_{1,0,-1,0}-240H_{1,0,0,0}-256H_{1,0,0,1}-64H_{1,0,1,1}-320H_{1,1,0,0}-64H_{1,1,0,1}-64H_{1,1,1,0}+64H_{1,1,1,1}+\frac{5408}{3}\right)z+\left(-\frac{19762}{27}\right)H_{0}+\left(-\frac{9620}{27}\right)H_{1}+\left(-\frac{928}{3}\right)H_{0,0,0}+\left(-\frac{640}{3}\right)H_{-1,-1,0}+\left(-\frac{1502}{9}\right)H_{0,0}+\left(-\frac{352}{3}\right)H_{-1,0,1}+\left(-\frac{566}{9}\right)H_{1,1}+\left(-\frac{434}{9}\right)H_{1,0}+\frac{220}{3}H_{1,1,1}+\frac{232}{3}H_{-1,0}+\frac{304}{3}\zeta_{3}+\frac{680}{3}H_{1,0,1}+\frac{680}{3}H_{1,1,0}+\frac{776}{3}H_{0,-1,0}+\frac{2390}{9}H_{0,1}+\frac{856}{3}H_{-1,0,0}+\frac{1240}{3}H_{1,0,0}+\frac{1342}{3}\zeta_{2}+\left(-\frac{224\zeta_{2}}{3}\right)H_{1}+\frac{32\zeta_{2}}{3}H_{-1}+\frac{856\zeta_{2}}{3}H_{0}+136H_{-1}\zeta_{3}+480H_{0}\zeta_{3}-248H_{1}\zeta_{3}+63\zeta_{4}-208\zeta_{2}H_{-1,-1}+112\zeta_{2}H_{-1,0}-88\zeta_{2}H_{0,-1}+316\zeta_{2}H_{0,0}-216\zeta_{2}H_{0,1}-48\zeta_{2}H_{1,0}+16\zeta_{2}H_{1,1}-40H_{0,0,1}+4H_{0,1,0}-56H_{0,1,1}-96H_{-1,-1,-1,0}+64H_{-1,-1,0,0}+160H_{-1,-1,0,1}+112H_{-1,0,-1,0}+8H_{-1,0,0,0}-64H_{-1,0,1,0}-32H_{-1,0,1,1}-176H_{0,-1,-1,0}+216H_{0,-1,0,0}+248H_{0,0,-1,0}-632H_{0,0,0,0}-388H_{0,0,0,1}-48H_{0,0,1,0}-96H_{0,0,1,1}+204H_{0,1,0,0}+144H_{0,1,0,1}+144H_{0,1,1,0}+128H_{0,1,1,1}+16H_{1,0,-1,0}+120H_{1,0,0,0}+128H_{1,0,0,1}+32H_{1,0,1,1}+160H_{1,1,0,0}+32H_{1,1,0,1}+32H_{1,1,1,0}-32H_{1,1,1,1}-\frac{4670}{3}\bigg]+N_{f}C_{F}^{2}\bigg[\left(48\zeta_{2}+48H_{-1,0}-48H_{0,0}\right)z^{2}+\left(-80\zeta_{2}H_{0}+80\zeta_{3}H_{0}+289H_{0}+600H_{1}-64H_{-1}\zeta_{2}-176H_{1}\zeta_{2}+612\zeta_{2}+992H_{1}\zeta_{3}-368\zeta_{3}+280\zeta_{4}-192\zeta_{2}H_{-1,0}+304H_{-1,0}+128\zeta_{2}H_{0,-1}-80\zeta_{2}H_{0,0}-350H_{0,0}+208\zeta_{2}H_{0,1}-340H_{0,1}-224\zeta_{2}H_{1,0}+592H_{1,0}-32\zeta_{2}H_{1,1}+32H_{1,1}-128H_{-1,-1,0}+352H_{-1,0,0}+56H_{0,0,0}+16H_{0,0,1}-16H_{0,1,0}+32H_{0,1,1}-64H_{1,0,0}-48H_{1,0,1}-144H_{1,1,0}-48H_{1,1,1}-384H_{-1,-1,0,0}-192H_{-1,0,-1,0}+352H_{-1,0,0,0}+256H_{0,-1,-1,0}+64H_{0,-1,0,0}-192H_{0,0,-1,0}+32H_{0,0,0,0}+16H_{0,0,0,1}+192H_{0,0,1,0}+128H_{0,0,1,1}-352H_{0,1,0,0}-208H_{0,1,0,1}-240H_{0,1,1,0}-112H_{0,1,1,1}-192H_{1,0,-1,0}+672H_{1,0,0,0}+160H_{1,0,0,1}+608H_{1,0,1,0}+128H_{1,0,1,1}+32H_{1,1,0,0}+160H_{1,1,0,1}+96H_{1,1,1,0}+64H_{1,1,1,1}+\frac{3747}{2}\right)z$
[…]